# Adapting Newton’s Method to Neural Networks through a Summary of Higher-Order Derivatives

### v3

@@ Appendix F Convergence rate in a simple case
$\mathbf{H}_{t}$ $:=\frac{\mathrm{d}^{2}\mathcal{L}}{\mathrm{d}\bm{\theta}^{2}}(\bm{\theta}_{t}),$ $\mathbf{g}_{t}$ $:=\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\bm{\theta}}(\bm{\theta}_{t}),$
@@
$\mathcal{L}(\bm{\theta}):=\frac{1}{2}\bm{\theta}^{T}\mathbf{H}\bm{\theta},$ (46)
@@
Theorem 1.
The method has a linear rate of convergence. For any $\bm{\theta}_{t}\neq 0$:
$\frac{\mathcal{L}(\bm{\theta}_{t+1})}{\mathcal{L}(\bm{\theta}_{t})}\leq\max_{s}\left(\frac{(A_{s}-a_{s})^{2}}{(A_{s}+a_{s})^{2}}\right),$
where $a_{s}=\min\mathrm{Sp}(\mathbf{H}_{s})$ and $A_{s}=\max\mathrm{Sp}(\mathbf{H}_{s})$. Moreover, this rate is optimal, since it is possible to build $\bm{\theta}_{t}$ such that:
$\frac{\mathcal{L}(\bm{\theta}_{t+1})}{\mathcal{L}(\bm{\theta}_{t})}=\max_{s}\left(\frac{(A_{s}-a_{s})^{2}}{(A_{s}+a_{s})^{2}}\right).$
@@
$\frac{\mathcal{L}(\bm{\theta}_{t+1})}{\mathcal{L}(\bm{\theta}_{t})}\leq\max_{s}\left(\frac{(\gamma_{s}-1)^{2}}{(\gamma_{s}+1)^{2}}\right),$
@@
F.1 Proof of Theorem 1
Proof.
@@ F.1 Proof of Theorem 1
$\mathcal{L}(\bm{\theta}_{t+1})$ $=\frac{1}{2}\bm{\theta}_{t+1}^{T}\mathbf{H}\bm{\theta}_{t+1}$
$=\frac{1}{2}(\bm{\theta}_{t}-\mathbf{G}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t})^{T}\mathbf{H}(\bm{\theta}_{t}-\mathbf{G}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t})$
$=\mathcal{L}(\bm{\theta}_{t})-\bm{\theta}_{t}^{T}\mathbf{H}\mathbf{G}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t}+\frac{1}{2}\bm{\eta}_{t}^{T}\mathbf{I}_{S:P}\mathbf{G}_{t}\mathbf{H}\mathbf{G}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t}$
$=\mathcal{L}(\bm{\theta}_{t})-\bm{\theta}_{t}^{T}\mathbf{H}\mathbf{G}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t}+\frac{1}{2}\bm{\eta}_{t}^{T}\bar{\mathbf{H}}_{t}\bm{\eta}_{t}$
$=\mathcal{L}(\bm{\theta}_{t})-\mathbf{g}_{t}^{T}\mathbf{G}_{t}\mathbf{I}_{P:S}\bar{\mathbf{H}}_{t}^{-1}\bar{\mathbf{g}}_{t}+\frac{1}{2}\bar{\mathbf{g}}_{t}\bar{\mathbf{H}}_{t}^{-1}\bar{\mathbf{g}}_{t}$
$=\mathcal{L}(\bm{\theta}_{t})-\frac{1}{2}\bar{\mathbf{g}}_{t}^{T}\bar{\mathbf{H}}_{t}^{-1}\bar{\mathbf{g}}_{t}.$
@@
$\leq-\min\left(\frac{2A_{s}a_{s}}{(A_{s}+a_{s})^{2}}\right)\bm{\theta}^{T}\mathbf{H}\bm{\theta}.$
Finally, when dividing by $\mathcal{L}(\bm{\theta}_{t})=\frac{1}{2}\bm{\theta}^{T}\mathbf{H}\bm{\theta}$, we have:
$\frac{\mathcal{L}(\bm{\theta}_{t+1})}{\mathcal{L}(\bm{\theta}_{t})}-1$ $\leq-\min\left(\frac{4A_{s}a_{s}}{(A_{s}+a_{s})^{2}}\right)$
$\frac{\mathcal{L}(\bm{\theta}_{t+1})}{\mathcal{L}(\bm{\theta}_{t})}$ $\leq\max\left(\frac{(A_{s}-a_{s})^{2}}{(A_{s}+a_{s})^{2}}\right)$
Besides, this rate is optimal, since it is possible to build $\bm{\theta}_{t}$ such that:
$\frac{\mathcal{L}(\bm{\theta}_{t+1})}{\mathcal{L}(\bm{\theta}_{t})}=\max_{s}\left(\frac{(A_{s}-a_{s})^{2}}{(A_{s}+a_{s})^{2}}\right).$
To do so, let $s_{0}\in\operatorname*{arg\,max}_{s}\left(\frac{(A_{s}-a_{s})^{2}}{(A_{s}+a_{s})^{2}}\right)$. Let $\mathbf{g}_{\min}$ be an eigenvector of $\mathbf{H}$ associated to $a_{s_{0}}$ and $\mathbf{g}_{\max}$ be an eigenvector of $\mathbf{H}$ associated to $A_{s_{0}}$, orthogonal with $\|\mathbf{g}_{\min}\|=\|\mathbf{g}_{\max}\|=1$. Also, let $\bm{\theta}_{t}=\mathbf{H}^{-1}(\mathbf{g}_{\min}+\mathbf{g}_{\max})$.
@@
$\mathcal{L}(\bm{\theta}_{t+1})-\mathcal{L}(\bm{\theta}_{t})=-\frac{1}{2}\frac{(\mathbf{g}_{s_{0}}^{T}\mathbf{g}_{s_{0}})^{2}}{\mathbf{g}_{s_{0}}^{T}\mathbf{H}_{s_{0}}\mathbf{g}_{s_{0}}}=-\frac{1}{2}\frac{2}{A_{s_{0}}+a_{s_{0}}}$
@@
$\frac{\mathcal{L}(\bm{\theta}_{t+1})}{\mathcal{L}(\bm{\theta}_{t})}$ $=1-\frac{1}{2}\frac{2}{A_{s_{0}}+a_{s_{0}}}\frac{1}{\frac{1}{2}\mathbf{g}_{s_{0}}^{T}\mathbf{H}_{s_{0}}^{-1}\mathbf{g}_{s_{0}}}$
@@
Appendix G Experimental details
@@
To implement the method proposed in Section 4, we propose Algorithm 4. The key function is compute_lr$(\lambda_{\mathrm{int}};\mathcal{L},\bm{\theta},\tilde{Z},\mathbf{u})$, which returns a solution $\bm{\eta}_{*}$ of:
$\bm{\eta}_{*}$ $=\left(\bar{\mathbf{H}}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\bm{\eta}_{*}\|\mathbf{D}^{2}\right)^{-1}\bar{\mathbf{g}},$
@@
$\quad\mathbf{D}$ $:=\mathrm{Diag}\left(\left(\left|\mathbf{D}^{(3)}_{\bm{\theta}}(\mathbf{u})\right|_{iii}^{1/3}\right)_{1\leq i\leq S}\right).$
“momentum$(\mu,\mathbf{x},\tilde{\mathbf{x}})$” returns $\mathbf{x}$ if $\tilde{\mathbf{x}}$ is undefined, else $\mu\tilde{\mathbf{x}}+(1-\mu)\mathbf{x}$. “schedule$(\tau_{\mathrm{sch}},p_{\mathrm{sch}},f_{\mathrm{sch}};\cdots)$” corresponds to torch.optim.lr_scheduler.ReduceLROnPlateau called every $\tau_{\mathrm{sch}}$ with patience $p_{\mathrm{sch}}$ and factor $f_{\mathrm{sch}}$, in order to reduce the step size $\lambda_{t}$ when the loss attains a plateau. [^6] See torch.optim.lr_scheduler.ReduceLROnPlateau. The samplers $\mathcal{D}_{g}$ and $\mathcal{D}_{\mathrm{newt}}$ are respectively used to compute the gradients $\mathbf{g}_{t}$ and $(\bar{\mathbf{H}},\bar{\mathbf{g}})$ used in “compute_lr”.
The hyperparameters are: the initial step size $\lambda_{1}$, the momentum $\mu_{g}$ on the gradients $\mathbf{g}_{t}$, the minibatch size $B$ to sample the $\tilde{Z}$ (used to compute $\bar{\mathbf{g}}$, $\bar{\mathbf{H}}$ and $\mathbf{D}$), the number of steps $\tau$ between each call of compute_lr, the momentum $\mu_{\eta}$ on the learning rates $\bm{\eta}_{t}$, the internal damping $\lambda_{\mathrm{int}}$, and the parameters of the scheduler $\tau_{\mathrm{sch}}$, $p_{\mathrm{sch}}$, $f_{\mathrm{sch}}$.
Algorithm 4 Complete implementation of the second-order optimization method described in Sec. 4.
@@
Hyperparams: $\lambda_{1},\mu_{g},B_{g},B,\tau,\mu_{\eta},\lambda_{\mathrm{int}},\tau_{\mathrm{sch}},p_{\mathrm{sch}},f_{\mathrm{sch}}$
@@
     $\mathcal{L}_{t}\leftarrow\mathcal{L}(\bm{\theta}_{t},Z_{t})$ (forward pass)
     $\mathbf{g}_{t}\leftarrow\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\bm{\theta}}(\bm{\theta}_{t},Z_{t})$ (backward pass)
@@
         $\bm{\eta}_{t}\leftarrow$ compute_lr$(\lambda_{\mathrm{int}};\mathcal{L},\bm{\theta}_{t},\tilde{Z}_{t},\tilde{\mathbf{g}}_{t})$
         $\tilde{\bm{\eta}}_{t}\leftarrow$ momentum$(\mu_{\eta};(\bm{\eta}_{t})_{+},\tilde{\bm{\eta}}_{t-1})$
@@
     $\bm{\theta}_{t+1}\leftarrow\bm{\theta}_{t}-\lambda_{t}\mathrm{Diag}(\tilde{\mathbf{g}}_{t})\mathbf{I}_{P:S}\tilde{\bm{\eta}}_{t}$ (training step)
     $\lambda_{t+1}\leftarrow$ schedule$(\tau_{\mathrm{sch}},p_{\mathrm{sch}},f_{\mathrm{sch}};t,\mathcal{L}_{t},\lambda_{t})$
@@
The “momentum” functions are used to deal with the stochastic part of the training process, since our method has not been designed to be robust against noise. The period $\tau$ is usually strictly greater than $1$, in order to avoid calling “compute_lr” at every step, which would be costly. The minibatch size $B$ should be large enough to reduce noise in the estimation of $\bm{\eta}_{*}$. If we denote by $B_{g}$ the size of the minibatches in $\mathcal{D}_{g}$, then we recommend the following setup: $\tau=\frac{B}{B_{g}}=\frac{1}{1-\mu_{g}}$. That way, we ensure that the training data are sampled from $\mathcal{D}_{g}$ and $\mathcal{D}_{\mathrm{newt}}$ at the same rate, and that $\tilde{\mathbf{g}}_{t}$ memorizes the preceding gradients $\mathbf{g}_{t}$ for $\tau$ steps. Besides, we have to take the positive part $(\bm{\eta}_{t})_{+}$ of $\bm{\eta}_{t}$ in order to avoid negative learning rates.
@@
$\mu_{g}$ $B_{g}$ $B$ $\tau$ $\mu_{\eta}$ $\tau_{\mathrm{sch}}$ $p_{\mathrm{sch}}$ $f_{\mathrm{sch}}$
$0.9$ $10^{2}$ $10^{3}$ $10$ $0.5$ $N_{e}$ $5$ $0.5$
@@
Ours $\lambda_{1}$ $10^{-1}$ $3\cdot 10^{-1}$ $3\cdot 10^{-1}$ $10^{-1}$
$\lambda_{\mathrm{int}}$ $10$ $3$ $3$ $10$
@@
Appendix H Very deep multilayer perceptron

### v4

@@ Appendix 0.F Convergence rate in a simple case
$\mathbf{H}_{t}$ $:=\frac{\mathrm{d}^{2}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{2}}(\boldsymbol{\theta}_{t}),$ $\mathbf{g}_{t}$ $:=\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}}(\boldsymbol{\theta}_{t}),$
@@
$\mathcal{L}(\boldsymbol{\theta}):=\frac{1}{2}\boldsymbol{\theta}^{T}\mathbf{H}\boldsymbol{\theta},$ (45)
@@
Theorem 0.F.1
The method has a linear rate of convergence. For any $\boldsymbol{\theta}_{t}\neq 0$:
$\frac{\mathcal{L}(\boldsymbol{\theta}_{t+1})}{\mathcal{L}(\boldsymbol{\theta}_{t})}\leq\max_{s}\left(\frac{(A_{s}-a_{s})^{2}}{(A_{s}+a_{s})^{2}}\right),$
where $a_{s}=\min\mathrm{Sp}(\mathbf{H}_{s})$ and $A_{s}=\max\mathrm{Sp}(\mathbf{H}_{s})$. Moreover, this rate is optimal, since it is possible to build $\boldsymbol{\theta}_{t}$ such that:
$\frac{\mathcal{L}(\boldsymbol{\theta}_{t+1})}{\mathcal{L}(\boldsymbol{\theta}_{t})}=\max_{s}\left(\frac{(A_{s}-a_{s})^{2}}{(A_{s}+a_{s})^{2}}\right).$
@@
$\frac{\mathcal{L}(\boldsymbol{\theta}_{t+1})}{\mathcal{L}(\boldsymbol{\theta}_{t})}\leq\max_{s}\left(\frac{(\gamma_{s}-1)^{2}}{(\gamma_{s}+1)^{2}}\right),$
@@
0.F.1 Proof of Theorem 0.F.1
Proof
@@
$\mathcal{L}(\boldsymbol{\theta}_{t+1})$ $=\frac{1}{2}\boldsymbol{\theta}_{t+1}^{T}\mathbf{H}\boldsymbol{\theta}_{t+1}$
$=\frac{1}{2}(\boldsymbol{\theta}_{t}-\mathbf{G}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t})^{T}\mathbf{H}(\boldsymbol{\theta}_{t}-\mathbf{G}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t})$
$=\mathcal{L}(\boldsymbol{\theta}_{t})-\boldsymbol{\theta}_{t}^{T}\mathbf{H}\mathbf{G}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t}+\frac{1}{2}\boldsymbol{\eta}_{t}^{T}\mathbf{I}_{S:P}\mathbf{G}_{t}\mathbf{H}\mathbf{G}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t}$
$=\mathcal{L}(\boldsymbol{\theta}_{t})-\boldsymbol{\theta}_{t}^{T}\mathbf{H}\mathbf{G}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t}+\frac{1}{2}\boldsymbol{\eta}_{t}^{T}\bar{\mathbf{H}}_{t}\boldsymbol{\eta}_{t}$
$=\mathcal{L}(\boldsymbol{\theta}_{t})-\mathbf{g}_{t}^{T}\mathbf{G}_{t}\mathbf{I}_{P:S}\bar{\mathbf{H}}_{t}^{-1}\bar{\mathbf{g}}_{t}+\frac{1}{2}\bar{\mathbf{g}}_{t}\bar{\mathbf{H}}_{t}^{-1}\bar{\mathbf{g}}_{t}$
$=\mathcal{L}(\boldsymbol{\theta}_{t})-\frac{1}{2}\bar{\mathbf{g}}_{t}^{T}\bar{\mathbf{H}}_{t}^{-1}\bar{\mathbf{g}}_{t}.$
@@
$\leq-\min\left(\frac{2A_{s}a_{s}}{(A_{s}+a_{s})^{2}}\right)\boldsymbol{\theta}^{T}\mathbf{H}\boldsymbol{\theta}.$
Finally, when dividing by $\mathcal{L}(\boldsymbol{\theta}_{t})=\frac{1}{2}\boldsymbol{\theta}^{T}\mathbf{H}\boldsymbol{\theta}$, we have:
$\frac{\mathcal{L}(\boldsymbol{\theta}_{t+1})}{\mathcal{L}(\boldsymbol{\theta}_{t})}-1$ $\leq-\min\left(\frac{4A_{s}a_{s}}{(A_{s}+a_{s})^{2}}\right)$
$\frac{\mathcal{L}(\boldsymbol{\theta}_{t+1})}{\mathcal{L}(\boldsymbol{\theta}_{t})}$ $\leq\max\left(\frac{(A_{s}-a_{s})^{2}}{(A_{s}+a_{s})^{2}}\right)$
Besides, this rate is optimal, since it is possible to build $\boldsymbol{\theta}_{t}$ such that:
$\frac{\mathcal{L}(\boldsymbol{\theta}_{t+1})}{\mathcal{L}(\boldsymbol{\theta}_{t})}=\max_{s}\left(\frac{(A_{s}-a_{s})^{2}}{(A_{s}+a_{s})^{2}}\right).$
To do so, let $s_{0}\in\operatorname*{arg\,max}_{s}\left(\frac{(A_{s}-a_{s})^{2}}{(A_{s}+a_{s})^{2}}\right)$. Let $\mathbf{g}_{\min}$ be an eigenvector of $\mathbf{H}$ associated to $a_{s_{0}}$ and $\mathbf{g}_{\max}$ be an eigenvector of $\mathbf{H}$ associated to $A_{s_{0}}$, orthogonal with $\|\mathbf{g}_{\min}\|=\|\mathbf{g}_{\max}\|=1$. Also, let $\boldsymbol{\theta}_{t}=\mathbf{H}^{-1}(\mathbf{g}_{\min}+\mathbf{g}_{\max})$.
@@
$\mathcal{L}(\boldsymbol{\theta}_{t+1})-\mathcal{L}(\boldsymbol{\theta}_{t})=-\frac{1}{2}\frac{(\mathbf{g}_{s_{0}}^{T}\mathbf{g}_{s_{0}})^{2}}{\mathbf{g}_{s_{0}}^{T}\mathbf{H}_{s_{0}}\mathbf{g}_{s_{0}}}=-\frac{1}{2}\frac{2}{A_{s_{0}}+a_{s_{0}}}$
@@
$\frac{\mathcal{L}(\boldsymbol{\theta}_{t+1})}{\mathcal{L}(\boldsymbol{\theta}_{t})}$ $=1-\frac{1}{2}\frac{2}{A_{s_{0}}+a_{s_{0}}}\frac{1}{\frac{1}{2}\mathbf{g}_{s_{0}}^{T}\mathbf{H}_{s_{0}}^{-1}\mathbf{g}_{s_{0}}}$
@@
Appendix 0.G Experimental details
@@
To implement the method proposed in Section 5, we propose Algorithm 2. The key function are compute_lr$(\lambda_{\mathrm{int}};\bar{\mathbf{H}},\bar{\mathbf{g}},\mathbf{D})$, which returns a solution $\boldsymbol{\eta}_{*}$ of:
$\boldsymbol{\eta}_{*}$ $=\left(\bar{\mathbf{H}}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\boldsymbol{\eta}_{*}\|\mathbf{D}^{2}\right)^{-1}\bar{\mathbf{g}},$
@@
$\quad\mathbf{D}$ $:=\mathrm{Diag}\left(\left(\left|\mathbf{D}^{(3)}_{\boldsymbol{\theta}}(\mathbf{u})\right|_{iii}^{1/3}\right)_{1\leq i\leq S}\right),$
and compute_Hg$(\mathcal{L},\boldsymbol{\theta},\tilde{Z},\mathbf{u})$, which returns the current value of $(\bar{\mathbf{H}},\bar{\mathbf{g}},\mathbf{D})$. “momentum$(\mu,\mathbf{x},\tilde{\mathbf{x}})$” returns $\mathbf{x}$ if $\tilde{\mathbf{x}}$ is undefined, else $\mu\tilde{\mathbf{x}}+\mathbf{x}$. “schedule$(p_{\mathrm{sch}},f_{\mathrm{sch}};\cdots)$” corresponds to torch.optim.lr_scheduler.ReduceLROnPlateau with patience $p_{\mathrm{sch}}$ and factor $f_{\mathrm{sch}}$, in order to reduce the step size $\lambda_{t}$ when the loss attains a plateau. [^6] See torch.optim.lr_scheduler.ReduceLROnPlateau. The samplers $\mathcal{D}_{g}$ and $\mathcal{D}_{\mathrm{newt}}$ are respectively used to compute the gradients $\mathbf{g}_{t}$ and $(\bar{\mathbf{H}},\bar{\mathbf{g}})$ used in “compute_lr”.
The hyperparameters are: the initial step size $\lambda_{1}$, the momentum $\mu_{g}$ on the gradients $\mathbf{g}_{t}$ (as for the SGD with momentum), the minibatch size $B$ to sample the $\tilde{Z}$ (used to compute $\bar{\mathbf{g}}$, $\bar{\mathbf{H}}$ and $\mathbf{D}$), the number of steps $\tau$ between each call of compute_lr, the averaging window $\upsilon$ for $\bar{\mathbf{H}},\bar{\mathbf{g}},\mathbf{D}$ (the average of $\bar{\mathbf{H}},\bar{\mathbf{g}},\mathbf{D}$ over the last $\tau\times\upsilon$ steps is used in compute_lr), the internal damping $\lambda_{\mathrm{int}}$, and the parameters of the scheduler $p_{\mathrm{sch}}$, $f_{\mathrm{sch}}$.
Algorithm 2 Implementation of the second-order optimization method described in Sec. 5.
@@
Hyperparams: $\lambda_{1},\mu_{g},B_{g},B,\tau,\upsilon,\lambda_{\mathrm{int}},p_{\mathrm{sch}},f_{\mathrm{sch}}$
@@
     $\mathcal{L}_{t}\leftarrow\mathcal{L}(\boldsymbol{\theta}_{t},Z_{t})$ (forward pass)
     $\mathbf{g}_{t}\leftarrow\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}}(\boldsymbol{\theta}_{t},Z_{t})$ (backward pass)
@@
         $\bar{\mathbf{H}}_{t},\bar{\mathbf{g}}_{t},\mathbf{D}_{t}\leftarrow$ compute_Hg$(\mathcal{L},\boldsymbol{\theta}_{t},\tilde{Z},\tilde{\mathbf{g}}_{t})$
         $\boldsymbol{\eta}_{t}\leftarrow$ compute_lr$(\lambda_{\mathrm{int}};\bar{\mathbf{H}},\bar{\mathbf{g}},\mathbf{D})$ ($\bar{\mathbf{H}},\bar{\mathbf{g}},\mathbf{D}$ averaged on the last $\upsilon$ steps)
         $\tilde{\boldsymbol{\eta}}_{t}\leftarrow$ momentum$(\mu_{\eta};(\boldsymbol{\eta}_{t})_{+},\tilde{\boldsymbol{\eta}}_{t-1})$
@@
     $\boldsymbol{\theta}_{t+1}\leftarrow\boldsymbol{\theta}_{t}-\lambda_{t}\mathrm{Diag}(\tilde{\mathbf{g}}_{t})\mathbf{I}_{P:S}\tilde{\boldsymbol{\eta}}_{t}$ (training step)
     $\lambda_{t+1}\leftarrow$ schedule$(p_{\mathrm{sch}},f_{\mathrm{sch}};t,\mathcal{L}_{t},\lambda_{t})$
@@
The “momentum” functions are used to deal with the stochastic part of the training process, since our method has not been designed to be robust against noise. The period $\tau$ is usually strictly greater than $1$, in order to avoid calling “compute_lr” at every step, which would be costly. The minibatch size $B$ should be large enough to reduce noise in the estimation of $\boldsymbol{\eta}_{*}$. If we denote by $B_{g}$ the size of the minibatches in $\mathcal{D}_{g}$, then we recommend the following setup: $\tau=\frac{B}{B_{g}}=\frac{1}{1-\mu_{g}}$. That way, we ensure that the training data are sampled from $\mathcal{D}_{g}$ and $\mathcal{D}_{\mathrm{newt}}$ at the same rate, and that $\tilde{\mathbf{g}}_{t}$ memorizes the preceding gradients $\mathbf{g}_{t}$ for $\tau$ steps. Besides, we have to take the positive part $(\boldsymbol{\eta}_{t})_{+}$ of $\boldsymbol{\eta}_{t}$ in order to avoid negative learning rates.
@@
$\mu_{g}$ $B_{g}$ $B$ $\tau$ $\upsilon$ $p_{\mathrm{sch}}$ $f_{\mathrm{sch}}$
$0.9$ $10^{2}$ $10^{3}$ $10$ $3$ $2$ $0.5$
@@
Ours $\lambda_{1}$ $3\cdot 10^{-2}$ $3\cdot 10^{-1}$ $3\cdot 10^{-1}$ $10^{-1}$
$\lambda_{\mathrm{int}}$ $1$ $1$ $1$ $3$
@@
Appendix 0.H Very deep multilayer perceptron